# A Dataset and Model for Realistic License Plate Deblurring

### v1

@@ 3.1 Data Collection
where ${v}$ denotes vehicular velocity, ${Et}$ represents exposure time, ${b}$ represents pixels blurred, ${D}$ is the distance between vehicle and camera, ${f}$ denotes camera focal length, and ${p}$ is pixel edge length on the sensor. Given that the actual speeds of individual vehicles are indeterminable, we standardized image captures on high-speed road sections with a regulatory speed limit of 70 km/h to ensure that ${D}$ remained within a range of 10-20 meters.
@@ 3.2 Data Post-processing
Due to the disparate ISO settings of the two cameras, Camera S and Camera L captured images with a different noise level. Consequently, during the conversion of RAW images to RGB format, wavelet denoising Liu et al. (2020) is employed after white-balancing, color mapping, and gamma correction.
@@ 5.4 Ablation Study
To evaluate the effectiveness of each proposed module, a series of ablation experiments are performed, which is shown in Table 3. The omission of the Latent Fusion Module led to a decline in global metrics, underscoring its effectiveness in fusing multi-scale features and improving the model’s performance in restoring sharp images. Removing the Text Reconstruction Module resulted in a significant downturn in global metrics, particularly noticeable under low light conditions. This highlights the pivotal role of the Text Reconstruction Module in enabling the model to have a deeper understanding and restoration capability for license plates affected by severe pixel disruption. Similarly, the exclusion of the Partition Discriminator Module deteriorated global metrics and notably affected the SSIM metric to a greater extent. This confirms the module’s contribution to enhancing the model focus on the details of each letter on the license plate.

### v2

@@ 3.1 Data Collection
where ${v}$ denotes vehicular velocity, ${Et}$ represents exposure time, ${b}$ represents pixels blurred, ${D}$ is the distance between vehicle and camera, ${f}$ denotes camera focal length, and ${p}$ is pixel edge length on the sensor. Given that the actual speeds of individual vehicles are indeterminable, we standardize image captures on high-speed road sections with a regulatory speed limit of 70 km/h to ensure that ${D}$ is remained within a range of 10-20 meters.
@@ 3.2 Data Post-processing
Due to the disparate ISO settings of the two cameras, Camera S and Camera L capture images with a different noise level. Consequently, during the conversion of RAW images to RGB format, wavelet denoising Liu et al. (2020) is employed after white-balancing, color mapping, and gamma correction.
@@ 5.4 Ablation Study
To evaluate the effectiveness of each proposed module, a series of ablation experiments are performed, which is shown in Table 3. The omission of the Latent Fusion Module leads to a decline in global metrics, underscoring its effectiveness in fusing multi-scale features and improving the model’s performance in restoring sharp images. Removing the Text Reconstruction Module results in a significant downturn in global metrics, particularly noticeable under low light conditions. This highlights the pivotal role of the Text Reconstruction Module in enabling the model to have a deeper understanding and restoration capability for license plates affected by severe pixel disruption. Similarly, the exclusion of the Partition Discriminator Module deteriorates global metrics and notably affects the SSIM metric to a greater extent. This confirms the module’s contribution to enhancing the model focus on the details of each letter on the license plate.